Chickentender37:

3 (x + y) = y If (x,y) is a solution to the equation above and y ≠ 0, what is the ratio of x/y

Tonycoolkid21:

surjithayer:

3(x+y)=y divide by y \[\frac{ 3(x+y) }{ y }=\frac{ y }{ y }\] \[\frac{ x+y }{ y }=\frac{ 1 }{ 3 }\] \[\frac{ x }{ y }+1=\frac{ 1 }{ 3 }\] \[\frac{ x }{ y }=\frac{ 1 }{ 3 }-1\] \[\frac{ x }{ y }=-\frac{ 2 }{ 3 }\]